# EHR-R1: A Reasoning-Enhanced Foundational Language Model for Electronic Health Record Analysis

### v1

@@ Abstract
Electronic Health Records (EHRs) contain rich yet complex information, and their automated analysis is critical for clinical decision-making. Despite recent advances of large language models (LLMs) in clinical workflows, their ability to analyze EHRs remains limited due to narrow task coverage and lack of EHR-oriented reasoning capabilities. This paper aims to bridge the gap, specifically, we present EHR-Ins, a large-scale, comprehensive EHR reasoning instruction dataset, comprising 300k high-quality reasoning cases and 4M non-reasoning cases across 42 distinct EHR tasks. Its core innovation is a thinking-graph-driven framework that enables to generate high-quality reasoning data at scale. Based on it, we develop EHR-R1, a series of reasoning-enhanced LLMs with up to 72B parameters tailored for EHR analysis. Through a multi-stage training paradigm, including domain adaptation, reasoning enhancement, and reinforcement learning, EHR-R1 systematically acquires domain knowledge and diverse reasoning capabilities, enabling accurate and robust EHR analysis. Lastly, we introduce EHR-Bench, a new benchmark curated from MIMIC-IV, spanning 42 tasks, to comprehensively assess reasoning and prediction across EHR scenarios. In experiments, we show that the resulting EHR-R1 consistently outperforms state-of-the-art commercial and open-source LLMs (including DeepSeek-V3 and GPT-4o), surpassing GPT-4o by over 30 points on MIMIC-Bench and achieving a 10% higher zero-shot AUROC on EHRSHOT. Collectively, EHR-Ins, EHR-R1, and EHR-Bench have significantly advanced the development for more reliable and clinically relevant EHR analysis.
@@ 1 INTRODUCTION
Electronic Health Records (EHRs) are comprehensive digital repositories of patient information, encompassing laboratory tests, medications, diagnoses, procedures, clinical notes, etc. [1, 2, 3, 4, 5]. Systematic analysis of EHRs is essential for modern healthcare, as accurate interpretation of patient histories enables early disease detection, personalized treatment planning, and improved clinical outcomes. In everyday clinical practice, physicians rely on electronic health records (EHRs) to address a wide spectrum of analytical needs, spanning decision-support queries, for example, “what is the most probable next diagnosis for this patient?”, to prognostic assessments, such as “what is the risk of hospital readmission within 60 days?” Advanced EHR analysis systems can substantially accelerate clinical workflows and augment medical decision-making, ultimately shaping the quality and timeliness of patient care.
Large language models (LLMs) have recently transformed biomedical natural language processing (BioNLP), achieving impressive results across a range of medical tasks [6, 7, 8, 9, 10, 11, 12, 13], yet their performance on EHR-related tasks remains limited [14, 15, 16, 17]. Even state-of-the-art commercial models struggle to extract, integrate, and reason over EHR data, which markedly limits their interoperability with daily used hospital information systems [18] and, consequently, their wide adoption in clinical practice.
Existing work on EHR analysis with LLMs is fundamentally constrained by two major challenges. First, in terms of task coverage, prior research has focused on narrow, task-specific objectives (e.g., risk prediction for a specific condition or outcome), typically restricted to particular diseases or event types, and has yet to deliver the holistic capabilities required to support evolving clinical workflows [19, 20, 21]. Second, in terms of reasoning ability, existing models struggle to construct reliable, EHR-oriented reasoning chains that demand both selective information filtering and multi-source integration. As a result, they are highly susceptible to the redundancy inherent in EHR data [22, 23] and struggle to integrate isolated findings into a coherent longitudinal understanding of disease progression [22, 24].
@@
We present EHR-Ins, a large-scale dataset capturing broad task diversity and explicit medical reasoning for EHR analysis. The corpus comprises 300K high-quality reasoning cases and 3.5 million non-reasoning cases spanning 42 EHR tasks. The tasks generally fall into two categories: decision-making (e.g., diagnosis and treatment recommendations) and risk-prediction (e.g., mortality and readmission). To ensure clinical fidelity and relevance, we develop a ‘thinking-graph’ driven reasoning data synthesis pipeline that (i) applies statistical analysis of entity co-occurrence ratios to identify key related entities, (ii) links them via knowledge from UMLS [25], and (iii) prompts GPT-4o to produce structured, step-by-step clinical reasoning. The resulting dataset is extensive and clinically grounded, enabling models to acquire diverse, context-rich reasoning capabilities.
@@
We evaluate EHR-R1 on EHRSHOT [19] and MIMIC-IV-CDM [26], as well as on a new benchmark, EHR-Bench, derived from MIMIC-IV. EHR-Bench spans 42 tasks across decision-making and risk-prediction settings, providing a balanced, comprehensive assessment of both reasoning and task-specific performance. Together, these benchmarks cover two distinct clinical centers—Stanford Medicine and Beth Israel Deaconess Medical Center—and encompass a broad range of EHR analysis tasks.
@@ 2.1 Training Data Overview
To enhance the LLM’s understanding of EHR data and its ability to identify relationships between medical entities within redundant EHR records, we propose a novel ‘thinking-graph’ pipeline to curate a new large-scale reasoning-enhanced EHR-analysis instruction data, EHR-Ins. This pipeline is designed to augment GPT-4o [27], enabling it to generate high-quality EHR reasoning chains. The details of this methodology can be found in Section 4.2.2. The synthesized reasoning data distribution for each analysis task in EHR-Ins is illustrated in Figure 2a, widely covering diverse decision-making tasks.
@@ 2.2.1 Benchmarks
To comprehensively evaluate LLM performance on EHR analysis, we introduce EHR-Bench, a benchmark derived from MIMIC-IV [28]. As summarized in Figure 3, EHR-Bench spans 12 subtypes and 42 tasks, organized into two major groups—decision making and risk prediction—covering a broad spectrum of clinically relevant settings [29, 30, 31].
@@
The MIMIC-IV-CDM [26] benchmark is also derived from MIMIC-IV. Compared with EHR-Bench, its distinct patient preprocessing pipeline and task design focus enable the evaluation of models’ generalization across domain shifts in EHR analysis tasks. Specifically, it targets diagnostic accuracy for four diseases—appendicitis, cholecystitis, diverticulitis, and pancreatitis—framed as decision-making tasks. We assess two diagnostic granularities: main disease diagnosis and ICD-level full diagnosis, enabling evaluation across coarse and fine levels of clinical specificity.
@@ 2.2.2 Metrics
Since large language models don’t have a traditional classification head, we get our probability estimates by using the tokens yes and no as our positive and negative classes, under a yes/no question prompt. At inference, we apply a technique called logit_biases to isolate the model’s scores for yes and no [6]. We then normalize these scores using a softmax function to get a probability for the positive class, which we use to calculate AUROC. Further metric details are provided in Section §4.4.2.
@@ 2.2.3 Baselines
  • Llama3.3-70B [32]: Developed by Meta, this series is a family of powerful open-source models that serve as a robust foundation for a wide range of applications.
  • Qwen2.5-72B [33]: As a versatile, general-purpose multilingual model from Alibaba Cloud, Qwen2.5 is recognized as one of the strongest open-source models available.
  • GPT-4o [27]: As OpenAI’s flagship closed-source model, GPT-4o is a powerful multimodal model known for its advanced reasoning and real-time responsiveness.
  • Qwen3-235B [34]: The reasoning model offers a unique ‘thinking’ and ‘non-thinking’ hybrid reasoning engine for enhanced problem-solving.
  • DeepSeek R1 [35]: This is the largest open-source reasoning model we evaluate, primarily focused on pushing the limits of reinforcement learning to achieve state-of-the-art results in complex reasoning.
  • GPT-OSS-120B [36]: Released by OpenAI, this open-source reasoning model provides strong reasoning and agentic capabilities to the open community.
  • Medgemma-27B [37]: Created by Google, this is a specialized medical model optimized for medical text and image comprehension, with a key multimodal capability.
  • OpenBioLLM-70B [38]: An open-source medical model meticulously fine-tuned for the biomedical field based on the Llama3-70B, consistently outperforming larger general-purpose models on biomedical benchmarks.
  • Baichuan-M2-32B [39]: This is a powerful open-source medical reasoning model based on Qwen3-32B. It has been specifically optimized for medical scenarios and trained from scratch on an unprecedented amount of high-quality medical data, enabling it to achieve deep medical expertise.
@@ 2.4 Generalization Evaluation on MIMIC-IV-CDM
The results of our generalization evaluation on the MIMIC-IV-CDM tasks are shown in Figure 6a. We focus on the zero-shot setting, where models are directly prompted [40] to perform EHR analysis tasks in MIMIC-IV-CDM without any task-specific training. The prompts used for evaluation are shown in Supplementary Table 5. We evaluate LLMs’ performance on two levels of diagnosis tasks: the main disease diagnoses and ICD coding level diagnoses. It can be observed that our model, EHR-R1-72B, achieves the highest performance on both types of tasks.
@@ 3 Discussion
Despite the remarkable progress of LLMs across diverse medical tasks, their application to clinical EHR analysis remains a significant challenge [14, 15, 16, 17]. Existing approaches fall short in two critical dimensions. First, task coverage remains narrow, as most efforts are confined to specific objectives or disease cohorts rather than enabling holistic analytical support for evolving clinical workflows. Second, reasoning ability is underdeveloped, with current LLMs struggling to generate reliable EHR-focused reasoning processes that progressively highlight key records, integrate fragmented evidence, and construct longitudinal, synthesized understanding of patient health.
@@
First comprehensive benchmark for EHR analysis. This paper introduces MIMIC-Bench, a new benchmark based on the MIMIC-IV dataset, which provides a comprehensive evaluation of LLMs on 42 diverse EHR tasks. This benchmark is designed to holistically assess a model’s ability to handle the full range of diverse clinical queries that exist in real-world EHRs, addressing coverage gaps in existing evaluations.
A robust and generalizable foundation model. We introduce EHR-R1, a state-of-the-art model that demonstrates superior performance on EHR analysis. By learning from the explicit reasoning paths in our data, our model not only sets new performance benchmarks on the proposed MIMIC-Bench but also shows exceptional zero-shot generalization capabilities on the out-of-distribution EHRSHOT dataset. This highlights EHR-R1’s robust ability to translate EHR content into clinical insight and understand temporal data to form a longitudinal narrative, which is critical for real-world clinical applications.
@@
Effective handling of diverse tasks. Trained on our EHR-Ins dataset, EHR-R1 performs a broad range of EHR tasks with high accuracy. Across 42 tasks in MIMIC-Bench—including clinical coding and other less-common or challenging tasks—it delivers strong, adaptive reasoning over diverse clinical queries. In addition, its zero-shot performance on unseen EHR data and tasks from the EHRSHOT dataset confirms that the model captures transferable structure rather than memorizing training data.
@@ 4.1 Problem Formulation
We formulate the EHR analysis problems in an instruction-tuning-based [41] generative framework.
@@
Figure 8: Overview of Data curation pipeline. The pipeline for our data curation process begins with the original MIMIC-IV dataset. From there, we establish EHR-Ins with reasoning enhanced EHR analysis instructions through thinking-graph–driven reasoning synthesis, along with a new comprehensive MIMIC-Bench.
@@ 4.2.1 MIMIC-IV Processing
Lastly, to convert the tabular observed EHR data $R_{\leq t}$ into a format that LLMs can process, we serialize the structured data[42, 43] into Markdown format[21], which is widely present in LLM pre-training corpora, facilitating better model comprehension. Specifically, each event is represented as plain text with a title (event name and standardized start time) and content. The content uses bullet-point syntax for events with single records and table syntax for others with multiple records, as shown in Case 1 and Case 2, respectively. The free-text case of EHR is shown in Case 5.
@@ 4.2.2 Thinking-Graph-Driven Reasoning Synthesis
Then we count the co-occurrence frequencies of the context entity $e_{i}\in\mathcal{E}_{\leq t}$ with the target label entities $y_{j}\in\mathcal{A}$ for a specific decision-making task denoted by its instruction $\mathcal{I}$ and calculate the Lift [44] between them. The Lift metric can discover strong associations within the data by measuring how much more frequently two entities appear together than expected by chance, allowing us to pinpoint the most relevant entities for the reasoning pathways. In our method, the Lift metric is calculated as:
@@
Although Lift can uncover potentially related medical entities from the historical EHR events, the medical relationships between these entities (including both the context entities in input and the label entities) and the labels remain uncovered and thus fail to capture the underlying clinical associations. To address this issue, we leverage the Unified Medical Language System (UMLS) [25] knowledge graph, which provides a structured representation of medical concepts and their relationships. UMLS integrates a wide range of medical terminologies and ontologies, including SNOMED CT, ICD, and MeSH, among others, making it an ideal resource to validate and strengthen the medical relationships between entities and labels in EHR data. Specifically, we map all medical entities identified in the EHR to their corresponding UMLS concepts. To increase the accuracy and completeness of this mapping, we use fuzzy matching provided in UMLS [25], allowing us to match medical terms in the EHR with UMLS concepts even when there are slight variations in spelling, abbreviations, or synonyms. This approach ensures a broader and more robust set of entity-to-concept mappings, improving the coverage of medical entities in the EHR data. For entities that cannot be mapped to UMLS concepts, we discard the related co-occurrence pairs, as these entities would not contribute valid medical relationships and could introduce noise into the analysis.

### v2

@@ Abstract
Electronic Health Records (EHRs) contain rich yet complex information, and their automated analysis is critical for clinical decision-making. Despite recent advances of large language models (LLMs) in clinical workflows, their ability to analyze EHRs remains limited due to narrow task coverage and lack of EHR-oriented reasoning capabilities. This paper aims to bridge the gap, specifically, we present EHR-Ins, a large-scale, comprehensive EHR reasoning instruction dataset, comprising 300k high-quality reasoning cases and 4M non-reasoning cases across 42 distinct EHR tasks. Its core innovation is a thinking-graph-driven framework that enables to generate high-quality reasoning data at scale. Based on it, we develop EHR-R1, a series of reasoning-enhanced LLMs with up to 72B parameters tailored for EHR analysis. Through a multi-stage training paradigm, including domain adaptation, reasoning enhancement, and reinforcement learning, EHR-R1 systematically acquires domain knowledge and diverse reasoning capabilities, enabling accurate and robust EHR analysis. Lastly, we introduce EHR-Bench, a new benchmark curated from MIMIC-IV, spanning 42 tasks, to comprehensively assess reasoning and prediction across EHR scenarios. In experiments, we show that the resulting EHR-R1 consistently outperforms state-of-the-art commercial and open-source LLMs (including DeepSeek-V3 and GPT-4o), surpassing GPT-4o by over 30 points on EHR-Bench and achieving a 10% higher zero-shot AUROC on EHRSHOT. Collectively, EHR-Ins, EHR-R1, and EHR-Bench have significantly advanced the development for more reliable and clinically relevant EHR analysis.
@@ 1 INTRODUCTION
Electronic Health Records (EHRs) are comprehensive digital repositories of patient information, encompassing laboratory tests, medications, diagnoses, procedures, clinical notes, etc. [sheikhalishahi2020benchmarking, DBLP:conf/nips/YecheKZHLFR21, DBLP:journals/jbi/SolaresRZRCTGPZ20, DBLP:conf/amia/XieZLTLRCCWDOGL22, DBLP:conf/chil/WangMCGHN20]. Systematic analysis of EHRs is essential for modern healthcare, as accurate interpretation of patient histories enables early disease detection, personalized treatment planning, and improved clinical outcomes. In everyday clinical practice, physicians rely on electronic health records (EHRs) to address a wide spectrum of analytical needs, spanning decision-support queries, for example, “what is the most probable next diagnosis for this patient?”, to prognostic assessments, such as “what is the risk of hospital readmission within 60 days?” Advanced EHR analysis systems can substantially accelerate clinical workflows and augment medical decision-making, ultimately shaping the quality and timeliness of patient care.
Large language models (LLMs) have recently transformed biomedical natural language processing (BioNLP), achieving impressive results across a range of medical tasks [nori2023capabilities, singhal2023large, liao2024automatic, qiu2025quantifying, DBLP:journals/npjdm/WuQLGLZWX25, DBLP:conf/acl/LiaoJW025, bedi2025medhelm, wu2025medreason], yet their performance on EHR-related tasks remains limited [rouhizadeh2025large, qu2024enhancing, soroush2024large, ren2025comprehensive]. Even state-of-the-art commercial models struggle to extract, integrate, and reason over EHR data, which markedly limits their interoperability with daily used hospital information systems [zhu2025healthflow] and, consequently, their wide adoption in clinical practice.
Existing work on EHR analysis with LLMs is fundamentally constrained by two major challenges. First, in terms of task coverage, prior research has focused on narrow, task-specific objectives (e.g., risk prediction for a specific condition or outcome), typically restricted to particular diseases or event types, and has yet to deliver the holistic capabilities required to support evolving clinical workflows [NEURIPS2023_d42db1f7, lin2025training, hegselmann2025large]. Second, in terms of reasoning ability, existing models struggle to construct reliable, EHR-oriented reasoning chains that demand both selective information filtering and multi-source integration. As a result, they are highly susceptible to the redundancy inherent in EHR data [wu2024instruction, chen2025narrative] and struggle to integrate isolated findings into a coherent longitudinal understanding of disease progression [wu2024instruction, kim2023detection].
@@
We present EHR-Ins, a large-scale dataset capturing broad task diversity and explicit medical reasoning for EHR analysis. The corpus comprises 300K high-quality reasoning cases and 3.5 million non-reasoning cases spanning 42 EHR tasks. The tasks generally fall into two categories: decision-making (e.g., diagnosis and treatment recommendations) and risk-prediction (e.g., mortality and readmission). To ensure clinical fidelity and relevance, we develop a ‘thinking-graph’ driven reasoning data synthesis pipeline that (i) applies statistical analysis of entity co-occurrence ratios to identify key related entities, (ii) links them via knowledge from UMLS [bodenreider2004unified], and (iii) prompts GPT-4o to produce structured, step-by-step clinical reasoning. The resulting dataset is extensive and clinically grounded, enabling models to acquire diverse, context-rich reasoning capabilities.
@@
We evaluate EHR-R1 on EHRSHOT [NEURIPS2023_d42db1f7] and MIMIC-IV-CDM [hager_evaluation_2024], as well as on a new benchmark, EHR-Bench, derived from MIMIC-IV. EHR-Bench spans 42 tasks across decision-making and risk-prediction settings, providing a balanced, comprehensive assessment of both reasoning and task-specific performance. Together, these benchmarks cover two distinct clinical centers—Stanford Medicine and Beth Israel Deaconess Medical Center—and encompass a broad range of EHR analysis tasks.
@@ 2.1 Training Data Overview
To enhance the LLM’s understanding of EHR data and its ability to identify relationships between medical entities within redundant EHR records, we propose a novel ‘thinking-graph’ pipeline to curate a new large-scale reasoning-enhanced EHR-analysis instruction data, EHR-Ins. This pipeline is designed to augment GPT-4o [hurst2024gpt], enabling it to generate high-quality EHR reasoning chains. The details of this methodology can be found in Section 4.2.2. The synthesized reasoning data distribution for each analysis task in EHR-Ins is illustrated in Figure 2a, widely covering diverse decision-making tasks.
@@ 2.2.1 Benchmarks
To comprehensively evaluate LLM performance on EHR analysis, we introduce EHR-Bench, a benchmark derived from MIMIC-IV [johnson2023mimic]. As summarized in Figure 3, EHR-Bench spans 12 subtypes and 42 tasks, organized into two major groups—decision making and risk prediction—covering a broad spectrum of clinically relevant settings [DBLP:journals/corr/abs-2502-06124, DBLP:journals/npjdm/RencJSWLBS24, DBLP:conf/ml4h/GuptaGCDPB22].
@@
The MIMIC-IV Clinical Decision Making (MIMIC-IV-CDM) [hager_evaluation_2024] benchmark is also derived from MIMIC-IV. Compared with EHR-Bench, its distinct patient preprocessing pipeline and task design focus enable the evaluation of models’ generalization across domain shifts in EHR analysis tasks. Specifically, it targets diagnostic accuracy for four diseases—appendicitis, cholecystitis, diverticulitis, and pancreatitis—framed as decision-making tasks. We assess two diagnostic granularities: main disease diagnosis and ICD-level full diagnosis, enabling evaluation across coarse and fine levels of clinical specificity.
@@ 2.2.2 Metrics
Since large language models don’t have a traditional classification head, we get our probability estimates by using the tokens yes and no as our positive and negative classes, under a yes/no question prompt. At inference, we apply a technique called logit_biases to isolate the model’s scores for yes and no [nori2023capabilities]. We then normalize these scores using a softmax function to get a probability for the positive class, which we use to calculate AUROC. Further metric details are provided in Section §4.4.2.
@@ 2.2.3 Baselines
  • Llama3.3-70B [DBLP:journals/corr/abs-2407-21783]: Developed by Meta, this series is a family of powerful open-source models that serve as a robust foundation for a wide range of applications.
  • Qwen2.5-72B [team2024qwen2]: As a versatile, general-purpose multilingual model from Alibaba Cloud, Qwen2.5 is recognized as one of the strongest open-source models available.
  • GPT-4o [hurst2024gpt]: As OpenAI’s flagship closed-source model, GPT-4o is a powerful multimodal model known for its advanced reasoning and real-time responsiveness.
  • Qwen3-235B [yang2025qwen3]: The reasoning model offers a unique ‘thinking’ and ‘non-thinking’ hybrid reasoning engine for enhanced problem-solving.
  • DeepSeek R1 [DBLP:journals/corr/abs-2501-12948]: This is the largest open-source reasoning model we evaluate, primarily focused on pushing the limits of reinforcement learning to achieve state-of-the-art results in complex reasoning.
  • GPT-OSS-120B [gptoss]: Released by OpenAI, this open-source reasoning model provides strong reasoning and agentic capabilities to the open community.
  • Medgemma-27B [sellergren2025medgemma]: Created by Google, this is a specialized medical model optimized for medical text and image comprehension, with a key multimodal capability.
  • OpenBioLLM-70B [dorfner2024biomedical]: An open-source medical model meticulously fine-tuned for the biomedical field based on the Llama3-70B, consistently outperforming larger general-purpose models on biomedical benchmarks.
  • Baichuan-M2-32B [wang2025baichuan]: This is a powerful open-source medical reasoning model based on Qwen3-32B. It has been specifically optimized for medical scenarios and trained from scratch on an unprecedented amount of high-quality medical data, enabling it to achieve deep medical expertise.
@@ 2.4 Generalization Evaluation on MIMIC-IV-CDM
The results of our generalization evaluation on the MIMIC-IV-CDM tasks are shown in Figure 6a. We focus on the zero-shot setting, where models are directly prompted [brown2020language] to perform EHR analysis tasks in MIMIC-IV-CDM without any task-specific training. The prompts used for evaluation are shown in Supplementary Table 5. We evaluate LLMs’ performance on two levels of diagnosis tasks: the main disease diagnoses and ICD coding level diagnoses. It can be observed that our model, EHR-R1-72B, achieves the highest performance on both types of tasks.
@@ 3 Discussion
Despite the remarkable progress of LLMs across diverse medical tasks, their application to clinical EHR analysis remains a significant challenge [rouhizadeh2025large, qu2024enhancing, soroush2024large, ren2025comprehensive]. Existing approaches fall short in two critical dimensions. First, task coverage remains narrow, as most efforts are confined to specific objectives or disease cohorts rather than enabling holistic analytical support for evolving clinical workflows. Second, reasoning ability is underdeveloped, with current LLMs struggling to generate reliable EHR-focused reasoning processes that progressively highlight key records, integrate fragmented evidence, and construct longitudinal, synthesized understanding of patient health.
@@
First comprehensive benchmark for EHR analysis. This paper introduces EHR-Bench, a new benchmark based on the MIMIC-IV dataset, which provides a comprehensive evaluation of LLMs on 42 diverse EHR tasks. This benchmark is designed to holistically assess a model’s ability to handle the full range of diverse clinical queries that exist in real-world EHRs, addressing coverage gaps in existing evaluations.
A robust and generalizable foundation model. We introduce EHR-R1, a state-of-the-art model that demonstrates superior performance on EHR analysis. By learning from the explicit reasoning paths in our data, our model not only sets new performance benchmarks on the proposed EHR-Bench but also shows exceptional zero-shot generalization capabilities on the out-of-distribution EHRSHOT dataset. This highlights EHR-R1’s robust ability to translate EHR content into clinical insight and understand temporal data to form a longitudinal narrative, which is critical for real-world clinical applications.
@@
Effective handling of diverse tasks. Trained on our EHR-Ins dataset, EHR-R1 performs a broad range of EHR tasks with high accuracy. Across 42 tasks in EHR-Bench—including clinical coding and other less-common or challenging tasks—it delivers strong, adaptive reasoning over diverse clinical queries. In addition, its zero-shot performance on unseen EHR data and tasks from the EHRSHOT dataset confirms that the model captures transferable structure rather than memorizing training data.
@@ 4.1 Problem Formulation
We formulate the EHR analysis problems in an instruction-tuning-based [ouyang2022training] generative framework.
@@
Figure 8: Overview of Data curation pipeline. The pipeline for our data curation process begins with the original MIMIC-IV dataset. From there, we establish EHR-Ins with reasoning-enhanced EHR analysis instructions through thinking-graph–driven reasoning synthesis, along with a new comprehensive EHR-Bench.
@@ 4.2.1 MIMIC-IV Processing
Lastly, to convert the tabular observed EHR data $R_{\leq t}$ into a format that LLMs can process, we serialize the structured data[liutapex, li2024table] into Markdown format[hegselmann2025large], which is widely present in LLM pre-training corpora, facilitating better model comprehension. Specifically, each event is represented as plain text with a title (event name and standardized start time) and content. The content uses bullet-point syntax for events with single records and table syntax for others with multiple records, as shown in Case 1 and Case 2, respectively. The free-text case of EHR is shown in Case 5.
@@ 4.2.2 Thinking-Graph-Driven Reasoning Synthesis
Then we count the co-occurrence frequencies of the context entity $e_{i}\in\mathcal{E}_{\leq t}$ with the target label entities $y_{j}\in\mathcal{A}$ for a specific decision-making task denoted by its instruction $\mathcal{I}$ and calculate the Lift [agrawal1993mining] between them. The Lift metric can discover strong associations within the data by measuring how much more frequently two entities appear together than expected by chance, allowing us to pinpoint the most relevant entities for the reasoning pathways. In our method, the Lift metric is calculated as:
@@
Although Lift can uncover potentially related medical entities from the historical EHR events, the medical relationships between these entities (including both the context entities in input and the label entities) and the labels remain uncovered and thus fail to capture the underlying clinical associations. To address this issue, we leverage the Unified Medical Language System (UMLS) [bodenreider2004unified] knowledge graph, which provides a structured representation of medical concepts and their relationships. UMLS integrates a wide range of medical terminologies and ontologies, including SNOMED CT, ICD, and MeSH, among others, making it an ideal resource to validate and strengthen the medical relationships between entities and labels in EHR data. Specifically, we map all medical entities identified in the EHR to their corresponding UMLS concepts. To increase the accuracy and completeness of this mapping, we use fuzzy matching provided in UMLS [bodenreider2004unified], allowing us to match medical terms in the EHR with UMLS concepts even when there are slight variations in spelling, abbreviations, or synonyms. This approach ensures a broader and more robust set of entity-to-concept mappings, improving the coverage of medical entities in the EHR data. For entities that cannot be mapped to UMLS concepts, we discard the related co-occurrence pairs, as these entities would not contribute valid medical relationships and could introduce noise into the analysis.